# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ III.2.2 Fluorographene
Fluorographene (FG) is a 2D stoichiometric graphene derivative (C1F1) material that exhibits a large band gap due to the out of plane carbon orbital saturated by bonding with fluorine. Some of the key properties of interest involve its large band gap, dielectric characteristics as well as potential surface physics and chemistry applications. Interestingly, the fundamental and optical band gaps as well as related derived properties has not been fully settled until recently. In particular, from experiments the onset of FG optical absorption spectrum has been estimated to lie between 3 eV - 5 eV. The fundamental gap from GW for the electron self-energy suggests a range of 7 eV to 8 eV. In addition, their excitonic effects appear to be very strong and the Bethe-Salpeter equation (BSE) confirmed the exciton binding energy of FG close to 2 eV, regardless of technicalities such as orbital sets or related parameters. In our study [203] we have employed both the highly accurate GW-BSE approach as well as DMC to ascertain these excitation characteristics. We found that the fundamental band gap from thoroughly converged GW and consistently extrapolated DMC methods agree within the error bars on value of 7.1(1) eV, establishing the reference for this material from two independent many-body approaches. Careful analysis of the BSE results confirmed 1.9 eV exciton binding and provided additional insights into the structure of the excitonic state based on the projection of the BSE exciton on the Bloch states.
III.2.3 Monolayer GeSe
@@
GeSe is a p-type semiconductor that has been widely studied because of its unique optical properties. Its bulk structure is generally known to possess a measured indirect gap of 1.07 eV to 1.29 eV at room temperature [214, 215, 216]. However, it was recently reported that a direct gap minimum of 1.3 eV was measured via optical spectroscopy [217]. This controversy is mainly due to the small difference between the direct and indirect gaps in bulk GeSe, and its high sensitivity in the method used for the measurement. In addition, DFT studies on GeSe showed strongly varied band gap energy and optimized lattice parameters depending on choice of XC functionals [218, 219, 220, 52]. These limitations on computing accurate optical and structure properties from DFT leads us to confirm that the use of more accurate method, which can describe exact optical and structural properties at the same time, is highly desired on the GeSe structure.
@@
We estimated the DMC direct gap of bulk GeSe to be 1.62(16) eV, which is in the excellent agreement with the experimental value of 1.53 eV [221]. Based on this DMC direct gap and our DMC quasiparticle gap result of 1.95(21) eV, bulk GeSe is expected to possess a weak exciton binding energy of 0.3 eV. Because the exact geometry for the GeSe monolayer is experimentally not known yet, its geometry is fully optimized using a surrogate Hessian-based parallel line search method [174] (similar to Ref. 109, 176). In the optimization process, the GeSe monolayer shows a shallow PES minimum over a large range of lattice parameters, which explains the high sensitivity of optimized DFT geometries to the choice of XC functional.
@@
In the PBE band structure for the GeSe monolayer, the DMC monolayer geometry shows a small direct gap of 1.50 eV at the point from the PBE result, which is competitive to the direct (1.24 eV) and indirect gaps (1.02 eV) located between the X and Y high symmetry point (see Fig. 9). Large differences in direct gap at the -point between the PBE (2.32 eV) and DMC geometries (1.50 eV) show that the monolayer band structure is very sensitive to its geometry, suggesting that controlling strain on the monolayer can be a route to manipulate the electronic and optical properties of monolayer GeSe. Additional details can be found in Ref. 175.
III.2.4 Monolayer GaSe and GaSxSe1-x Alloys
@@
Post Transition Metal Chalcogenides (PTMCs) are a class of 2D materials that have suitable band gaps for photovoltaics and transistors[222, 223, 224, 225, 226, 227, 228, 229, 230, 231] and have lower exciton binding energies than TMDs [222, 223, 224], which make them suitable for water-splitting applications [232]. Monolayer GaSe is a PTMC that has been reliably synthesized and measurements of the quasiparticle gap, optical gap, and lattice constant have been performed [43, 44, 45, 46, 47, 48, 49, 41, 50]. The experimental lattice constant of monolayer GaSe has a measured value of Å[47]. In addition, 2D GaSe has an experimental indirect band gap of 3.5 eV (on a graphene substrate) [43], which is much larger than the bulk GaSe band gap of 2.0 eV. The optical band gap has been measured to be 3.3 eV [45], which implies an exciton binding energy lower than 0.2 eV. Despite these well-characterized measurements, computational results can significantly vary based on which functional is used [204].
@@
The most apparent discrepancy is the location of the conduction band edge at each high symmetry k-point. The energy differences are so small between each high symmetry point ( 0.2 eV to 0.3 eV) that different functionals/pseudopotentials can predict the gap value to have different values and some methods even incorrectly predict the gap to be direct. Improvements to the underestimation of the gap can be achieved using methods such as GW or BSE, but these results depend on which functional is used to generate the starting wavefunction and the indirect/direct discrepancy still exists. We used DMC to obtain the optimal lattice constant by isotropically scaling the lattice and finding the energy minima. To demonstrate the weak dependence that DMC has on the starting functional, we created DMC the trial wavefunction with PBE, LDA and SCAN functionals and performed the same calculation. We computed 3.74(2) Å for DMC-PBE, 3.75(1) Å for DMC-LDA, and 3.75(1) Å for DMC-SCAN, which is in close agreement with experiment. With DMC, we confirmed that 2D GaSe is an indirect material (-M) with a quasiparticle gap of 3.69(5) eV, which is in excellent agreement with the experimental value. From our calculation of the optical band gap, we obtained a maximum bound on the exciton binding energy to be 80 meV, which confirms that GaSe is a 2D material with a lower exciton binding energy than most TMDs [9], making it suitable for water-splitting applications. More details of this work can be found in Ref. 204.
@@
We extended our approach to 2D GaSxSe1-x alloys [237], since alloying is a promising technique to control the properties of single and few layer structures [238, 239, 240, 241, 242, 243, 244, 245], which includes PTMCs[246, 247, 248, 41, 249]. Studying alloys with accurate computational methods is a difficult endeavor because usually the local atomic ordering and stoichiometry of the different alloyed structures is unknown. In order to tackle this issue, ab-initio methods have been coupled with methods such as Cluster Expansion to construct the energy hull diagram of an alloy system, which alloys us to determine if a structure is thermodynamically stable and possibly exist in nature [250].
@@
In this work, we designed a high-throughput workflow to compute the energy hull diagram of an alloy system with QMC methods, using 2D GaSxSe1-x as a case study. To achieve this, we proposed a method we deem Jastrow sharing, which involves recycling the optimized Jastrow parameters between various alloys with different stoichiometries. Specifically, we optimized the Jastrow parameters of the alloy with the smallest Wigner-Seitz radius and used those parameters for other structures of interest in the alloy system. We demonstrated that this eliminates the need for unnecessary VMC Jastrow optimization simulations and can reduce the total computational time by 1/4 (75 of the original computational time). After testing the validity of this Jastrow sharing approach, we went on to compute the alloy formation energies with DMC (extrapolated to the thermodynamic limit) for selected points on the energy hull diagram. This method can easily be implemented for other 2D alloy systems where the Jastrow sensitivities of the pseudopotentials of the atoms in the system are low, which can eventually aid in accurate studies of more complex alloy systems (i.e. alloying 2D transition metal oxide materials). Additional information from this study can be found in Ref. 237.
III.2.5 Spin-orbit effects, topological states
@@
Layered RuCl3 for Kitaev’s spin liquid: Recently, the QMC method has been extended to spin-orbit Hamiltonians with trial functions built upon two-component spinors and the spin-space sampled from an overcomplete set of spin configurations [251, 252]. For the first time we have applied this methodology to layered RuCl3 [122], where claims of sizeable contributions of Ru atomic spin-orbit to gap opening have been claimed [253, 254]. Our calculations did not find support for this scenario. We found that the quasi-particle band gap opens due to expected significant Hubbard repulsion on the Ru atom. That proved to be the dominant effect while spin-orbit provides a mild shift of approximately 0.2 eV, as also expected for the intermediate strength of the atomic spin-orbit for the Ru atom (see Fig.12).
@@
Somewhat larger impact of the spin-orbit was observed in the cohesive energy, however, this is caused almost exclusively by the larger shift in high symmetry isolated Ru atom, not in the solid where the hybridization partially quenches and averages the corresponding splittings. We note that RuCl3 is considered to be a promising candidate for realization of Kitaev spin liquid with low-lying collective states. However, that these are in the meV range (i.e., much smaller than the difference between ferromagnetic and antiferromagnetic states), hence outside the statistical resolution of current valence electronic structure QMC. Even if such wavefunction could be constructed, the presence of Kitaev physics presence would be hidden in the statistical noise and very extensive calculations would be needed to discern such a signal. Additional information from this work can be found in Ref. 122.
@@
Layered TbMn6Sn6 for Chern magnets: A distinct class of materials, RMn6Sn6, where R denotes a rare-earth element, displays rich and intricate physical phenomena such as strong electron-electron correlations, spin-orbit effects as well as possibilities of forming states with topological order.This stems from structural peculiarities with 2D Kagome layers of Mn interlaced by R and Sn layers. In this respect, the rare-earth Tb atom is of particular interest since it is the only R-element that forms an out-of-plane spin order compound. Experiments showed indications that TbMn6Sn6 is close to realizing a quantum-limit Chern magnet, as predicted by the Haldane model. Indeed, Kagome lattice geometry with an out-of-plane magnetization formed by the Mn atoms and the presence of strong spin-orbit coupling (SOC) originating from the Tb and Sn atoms provide the necessary conditions for opening the Chern gap. In our very recent study Ref.255, we use DMC and DFT with Hubbard U (DFT+U) calculations to examine the electronic structure of TbMn6Sn6. We find that DFT+U and single-reference QMC calculations exhibit the same overestimation of the magnetic moments as meta-GGA and hybrid density functional approximations. Our findings point to the need for improved orbitals as well as wavefunctions for this class of materials. We have concluded that significant multi-reference effects have to be included in order to capture the static correlations which are necessary for an accurate prediction of magnetic properties. We have probed for the occurrence of topological order, where we have explored DFT+U with Mn magnetic moments adjusted to experiment. We have observed the Dirac crossing in bulk to be close to the Fermi level, within 120 meV, in agreement with the experiments. The possibility of crossing the Fermi level has been further enhanced in non-stoichiometric slab calculations, keeping realization of Chern magnetism in this limit within the experimental reach. More information can be found in Ref. 255.
@@ III.3 Interlayer Interactions
Layering 2D materials offers additional degrees of freedom which allow us to explore the correlation between interlayer coupling and physical properties. This exploration can lead to the emergence of novel electronic states in 2D materials. Interlayer coupling in these layered 2D materials varies widely, ranging from purely dispersive interactions, such as those seen in bilayer graphene [256, 257], to complex interactions involving weak interlayer chemical bonding and even metallic bonding, in addition to vdW interactions. This complexity in interlayer interaction has been verified through X-ray density investigations for a 2D transition metal dichalcogenide material [258]. It is important to note that vdW interactions among these interactions are a many-body phenomenon stemming from non-local electronic correlations induced by the instantaneous fluctuation of electron density. These interactions may not be adequately captured by standard DFT frameworks, highlighting the need for a more sophisticated theoretical framework that can address these interactions equally. Such a framework is essential for capturing the diverse interlayer coupling which could vary depending on different stacking modes or twisting degrees of freedom. To illustrate this point, we present examples of DMC calculations aimed at investigating the nature of interlayer binding in 2D layered systems, providing a comprehensive understanding of interlayer coupling beyond what DFT offers.

### v2

@@ III.2.2 Fluorographene
Fluorographene (FG) is a 2D stoichiometric graphene derivative (C1F1) material that exhibits a large band gap due to the complete out-of-plane bonding of carbon orbitals with fluorine (C-F sp3 bonds). Some of the key properties of interest involve its large band gap, dielectric characteristics, and potential surface physics and chemistry applications. Interestingly, the fundamental and optical band gaps as well as related derived properties have not been fully settled until recently. In particular, from experiments, the onset of FG optical absorption spectrum has been estimated to lie between 3 eV and 5 eV. The fundamental gap from GW for the electron self-energy suggests a range of 7 eV to 8 eV. In addition, their excitonic effects appear to be very strong, and the BSE confirmed the exciton binding energy of FG close to 2 eV, regardless of technicalities such as orbital sets or related parameters. In our study, [173] we have employed both the highly accurate GW-BSE approach and DMC to ascertain these excitation characteristics. We found that the fundamental band gap from thoroughly converged GW and consistently extrapolated DMC methods agree within the error bars on value of 7.1(1) eV, establishing the reference for this material from two independent many-body approaches. Careful analysis of the BSE results confirmed 1.9 eV exciton binding and provided additional insights into the structure of the excitonic state based on the projection of the BSE exciton on the Bloch states.
III.2.3 Monolayer hBN
Monolayer hBN is a promising insulating 2D material with potential applications in electronic devices such as vertical tunneling diodes [240, 241] and supercapacitors.[242] In addition, hBN is an ideal substrate material for graphene-based electronics due to its similar lattice constant.[243, 244, 245, 246] Recently, Hunt et al. [148] studied monolayer hBN through the lens of many-body techniques (DMC and GW). The DMC quasiparticle gap of monolayer hBN was found to be 8.8(3) eV and indirect (from - ). This value was significantly higher than those predicted by G0W0 and GW0 (ranging from 7.43 eV to 7.72 eV). Fig. 8 depicts the FS convergence of the DMC quasiparticle and excitonic gaps of monolayer hBN. It was demonstrated that the DMC quasiparticle gap falls off as a function of but can be corrected by subtracting a screened Madelung constant [247] from the gap. From DFT calculations, a large 0.73 eV vibrational correction was found for monolayer hBN. With regard to excitonic effects, an exciton binding energy of 1.9(4) eV was found for the indirect exciton ( - ), and 1.8(4) eV was found for the direct ( - ), where the results for the direct exciton are in excellent agreement with the GW-BSE results performed in this study. A possible major source of error for this system (overestimation of the gap) could be unaccounted fixed-node errors from using a single-determinant wavefunction. For example, it is entirely possible that the fixed-node error does not equally cancel when computing energy differences for the ground and excited states. It is also possible that the vibrational renormalization of the gap is underestimated. Further information can be found in Ref. 148.
III.2.4 Monolayer GeSe
@@
GeSe is a p-type semiconductor that has been widely studied because of its unique optical properties. Its bulk structure is generally known to possess a measured indirect gap of 1.07 eV to 1.29 eV at room temperature.[248, 249, 250] However, it was recently reported that a direct gap minimum of 1.3 eV was measured via optical spectroscopy.[251] This controversy is mainly due to the small difference between the direct and indirect gaps in bulk GeSe, and its high sensitivity in the method used for the measurement. In addition, DFT studies on GeSe showed strongly varied band gap energy and optimized lattice parameters depending on choice of XC functionals.[252, 253, 254, 52] These limitations on computing accurate optical and structure properties from DFT lead us to confirm that the use of a more accurate method, which can describe exact optical and structural properties at the same time, is highly desired on the GeSe structure.
@@
We estimated the DMC direct gap of bulk GeSe to be 1.62(16) eV, which is in excellent agreement with the experimental value of 1.53 eV.[255] Based on this DMC direct gap and our DMC quasiparticle gap result of 1.95(21) eV, bulk GeSe is expected to possess a weak exciton binding energy of 0.3 eV. Because the exact geometry for the GeSe monolayer is experimentally not known yet, its geometry is fully optimized using a surrogate Hessian-based parallel line-search method [209] (similar to Ref. 109, 174). In the optimization process, the GeSe monolayer shows a shallow PES minimum over a large range of lattice parameters, which explains the high sensitivity of optimized DFT geometries to the choice of XC functional.
@@
In the PBE band structure for the GeSe monolayer, the DMC monolayer geometry shows a small direct gap of 1.50 eV at the point from the PBE result, which is competitive to the direct (1.24 eV) and indirect gaps (1.02 eV) located between the X and Y high symmetry point (see Fig. 9). Large differences in the direct gap at the -point between the PBE (2.32 eV) and DMC geometries (1.50 eV) show that the monolayer band structure is very sensitive to its geometry, suggesting that controlling strain on the monolayer can be a route to manipulate the electronic and optical properties of monolayer GeSe. Additional details can be found in Ref. 146.
III.2.5 Monolayer GaSe and GaSxSe1-x Alloys
@@
PTMCs are a class of 2D materials that have suitable band gaps for photovoltaics and transistors[256, 257, 258, 259, 260, 261, 262, 263, 264, 265] and have lower exciton binding energies than TMDs,[256, 257, 258] which make them suitable for water-splitting applications.[266] Monolayer GaSe is a PTMC that has been reliably synthesized, and measurements of the quasiparticle gap, optical gap, and lattice constant have been performed.[43, 44, 45, 46, 47, 48, 49, 41, 50] The experimental lattice constant of monolayer GaSe has a measured value of Å.[47] In addition, 2D GaSe has an experimental indirect band gap of 3.5 eV (on a graphene substrate),[43] which is much larger than the bulk GaSe band gap of 2.0 eV. The optical band gap has been measured to be 3.3 eV,[45] which implies an exciton binding energy lower than 0.2 eV. Despite these well-characterized measurements, computational results can significantly vary based on which functional is used.[232]
@@
The most apparent discrepancy is the location of the conduction band edge at each high symmetry -point. The energy differences are so small between each high symmetry point (0.2 eV to 0.3 eV) that different functionals/pseudopotentials can predict the gap value to have different values, and some methods even incorrectly predict the gap to be direct. Improvements to the underestimation of the gap can be achieved using methods such as GW or BSE, but these results depend on which functional is used to generate the starting wavefunction, and the indirect/direct discrepancy still exists. We used DMC to obtain the optimal lattice constant by isotropically scaling the lattice and finding the energy minima. To demonstrate the weak dependence that DMC has on the starting functional, we created the DMCtrial wavefunction with PBE, LDA, and SCAN functionals and performed the same calculation. We computed 3.74(2) Å for DMC-PBE, 3.75(1) Å for DMC-LDA, and 3.75(1) Å for DMC-SCAN, which is in close agreement with experiment. With DMC, we confirmed that 2D GaSe is an indirect material (-M) with a quasiparticle gap of 3.69(5) eV, which is in excellent agreement with the experimental value. A full comparison of the electronic properties of 2D GaSe computed with DFT, G0W0 and DMC is depicted in Fig. 10. From our calculation of the optical band gap, we obtained a maximum bound on the exciton binding energy to be 80 meV, which confirms that GaSe is a 2D material with a lower exciton binding energy than most TMDs,[9] making it suitable for water-splitting applications. More details of this work can be found in Ref. 232.
@@
We extended our approach to 2D GaSxSe1-x alloys [271] since alloying is a promising technique to control the properties of single- and few-layer structures,[272, 273, 274, 275, 276, 277, 278, 279] which includes PTMCs.[280, 281, 282, 41, 283] Studying alloys with accurate computational methods is a difficult endeavor because usually the local atomic ordering and stoichiometry of the different alloyed structures are unknown. To tackle this issue, ab initio methods have been coupled with methods such as cluster expansion to construct the energy hull diagram of an alloy system, which allows us to determine whether a structure is thermodynamically stable and possibly can exist in nature.[284]
@@
In this work, we designed a high-throughput workflow to compute the energy hull diagram of an alloy system with QMC methods, using 2D GaSxSe1-x as a case study. To achieve this, we proposed a method we call Jastrow sharing, which involves recycling the optimized Jastrow parameters between various alloys with different stoichiometries. Specifically, we optimized the Jastrow parameters of the alloy with the smallest Wigner-Seitz radius and used those parameters for other structures of interest in the alloy system. We demonstrated that this eliminates the need for unnecessary VMC Jastrow optimization simulations and can reduce the total computational time by 1/4. After testing the validity of this Jastrow sharing approach, we went on to compute the alloy formation energies with DMC (extrapolated to the thermodynamic limit) for selected points on the energy hull diagram (see Fig. 11). This method can easily be implemented for other 2D alloy systems where the Jastrow sensitivities of the pseudopotentials of the atoms in the system are low, which can eventually aid in accurate studies of more complex alloy systems (i.e., alloying 2D transition metal oxide materials). Additional information from this study can be found in Ref. 271.
III.2.6 Spin-Orbit Effects, Topological States
@@
Layered RuCl3 for Kitaev’s spin liquid: Recently, the QMC method has been extended to spin-orbit Hamiltonians with trial functions built upon two-component spinors and the spin-space sampled from an overcomplete set of spin configurations.[136, 176] For the first time, we have applied this methodology to layered RuCl3,[118] where claims of sizeable contributions of the Ru atomic spin-orbit to the gap opening have been made.[285, 286] Our calculations did not find support for this scenario. We found that the quasiparticle band gap opens due to expected significant Hubbard repulsion on the Ru atom. That proved to be the dominant effect, while spin-orbit provides a mild shift of approximately 0.2 eV, as also expected for the intermediate strength of the atomic spin-orbit for the Ru atom (see Fig. 12).
@@
A somewhat larger impact of the spin-orbit was observed in the cohesive energy; however, this is caused almost exclusively by the larger shift in the high symmetry isolated Ru atom, not in the solid where the hybridization partially quenches and averages the corresponding splittings. We note that RuCl3 is considered to be a promising candidate for realization of Kitaev spin liquid with low-lying collective states. However, these are in the meV range (i.e., much smaller than the difference between FM and AFM states), hence outside the statistical resolution of current valence electronic structure QMC. Even if such wavefunction could be constructed, the presence of Kitaev physics would be hidden in the statistical noise, and very extensive calculations would be needed to discern such a signal. Additional information from this work can be found in Ref. 118.
@@
Layered TbMn6Sn6 for Chern magnets: A distinct class of materials, RMn6Sn6, where R denotes a rare-earth element, displays rich and intricate physical phenomena such as strong electron-electron correlations, spin-orbit effects, and possibilities of forming states with topological order. This stems from structural peculiarities with 2D Kagome layers of Mn interlaced by R and Sn layers. In this respect, the rare-earth Tb atom is of particular interest since it is the only R element that forms an out-of-plane spin order compound. Experiments indicated that TbMn6Sn6 is close to realizing a quantum-limit Chern magnet, as predicted by the Haldane model. Indeed, Kagome lattice geometry with an out-of-plane magnetization formed by the Mn atoms and the presence of strong SOC originating from the Tb and Sn atoms provide the necessary conditions for opening the Chern gap. In our very recent study (Ref. 147), we use DMC and DFT with Hubbard U (DFT+U) calculations to examine the electronic structure of TbMn6Sn6. We find that DFT+U and single-reference QMC calculations exhibit the same overestimation of the magnetic moments as meta-GGA and hybrid density functional approximations. Our findings point to the need for improved orbital quality and for potentially extending beyond single-determinant wavefunctions for this class of materials. We have concluded that significant multireference effects must be included to capture the static correlations necessary for an accurate prediction of magnetic properties. We have probed for the occurrence of topological order, where we have explored DFT+U with Mn magnetic moments adjusted to experiment. We have observed the Dirac crossing in bulk to be close to the Fermi level, within 120 meV, in agreement with the experiments. The possibility of crossing the Fermi level has been further enhanced in nonstoichiometric slab calculations, keeping realization of Chern magnetism in this limit within the experimental reach. More information can be found in Ref. 147.
Enhancing MnBi2Te4 stability: MnBi2Te4 has an ordered layer of Mn atoms that lead to the observation of the quantum anomalous Hall effect when all the magnetic moments of its layers are aligned ferromagnetically.[287] However, it is shown that departure from the crystalline long-range order can lead to contradictory observations with regard to the existence of a Dirac cone or a gapless topological surface state.[288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302] Several studies show that Mn atoms migrating to Bi sites and creating MnBi defects can create a ferrimagnetic configuration and strongly modify the topological ground state.[301, 303, 304, 305] In our recent study (Ref. 306), we evaluate the potential for increasing the Mn-Bi antisite formation through chemical doping. Using a comprehensive screening set using DFT and DMC calculations, we find that group III elements, such as Sc, Y, and La dopants on Bi sites, have the potential to increase the formation energy of MnBi defects. None the dopants we studied on Mn sites has increased the MnBi defect formation energy. We show that the low antisite formation energy of MnBi defects in pristine MnBi2Te4 originates from internal strain between the Bi-Te and Mn-Te layers. Group III dopants on Bi sites decrease the coupling between these two layers by increasing the bond lengths between the Bi atoms and the Te atoms that are shared between the Bi and Te atoms. More information can be found in Ref. 306.
@@ III.3 Interlayer Interactions
Layering 2D materials offers additional degrees of freedom, thereby enabling us to explore the correlation between interlayer coupling and physical properties. This exploration can lead to the emergence of novel electronic states in 2D materials. Interlayer coupling in these layered 2D materials varies widely, ranging from purely dispersive interactions, such as those seen in bilayer graphene,[307, 308] to complex interactions involving weak interlayer chemical bonding and even metallic bonding, in addition to vdW interactions. This complexity in interlayer interaction has been verified through x-ray density investigations for a 2D TMD material.[309] It is important to note that vdW interactions among these interactions are a many-body phenomenon stemming from nonlocal electronic correlations induced by the instantaneous fluctuation of electron density. These interactions may not be adequately captured by standard DFT frameworks, highlighting the need for a more sophisticated theoretical framework that can address these interactions equally. Such a framework is essential for capturing the diverse interlayer coupling, which could vary depending on different stacking modes or twisting degrees of freedom. To illustrate this point, we present examples of DMC calculations aimed at investigating the nature of interlayer binding in 2D layered systems, providing a comprehensive understanding of interlayer coupling beyond what DFT offers.